# From the Least to the Most: Building a Plug-and-Play Visual Reasoner via Data Synthesis

### v1

@@ Abstract
Chuanqi Cheng1,2  Jian Guan2 thanks:   Equal Contribution.  Wei Wu2 thanks:   Corresponding Author.  Rui Yan1 1Gaoling School of Artificial Intelligence, Renmin University of China  2Ant Group {chengchuanqi, ruiyan}@ruc.edu.cn {jianguanthu, wuwei19850318}@gmail.com
@@ 1 Introduction
In this work, we explore vision language models (VLMs) as a typical example of LMMs. Despite the rapid progress, state-of-the-art VLMs still face challenges in reasoning over visual content. As exemplified in Figure 1, intuitively, the question can be correctly solved following a “least-to-most” paradigm Zhou et al. (2022), in which the question is decomposed into a series of sub-questions, and an answer is deduced by resolving the sub-questions step by step. The problem is that existing VLMs are inept at performing such multi-step reasoning because (1) Multi-step reasoning paths, like the one shown in Figure 1 (right), are rarely included in the training data Dai et al. (2024) [^1] A few datasets, such as A-OKVQA Schwenk et al. (2022), VCR Zellers et al. (2019), and ScienceQA Lu et al. (2022), contain rationales. However, the rationales merely provide explanations for the answers and thus differ significantly from the multi-step reasoning data we study in the work.. The VLMs have few opportunities to develop the reasoning capability from the subsequent post-training. (2) Different from reasoning over text, solving questions in a vision-language context may require manipulating the input image (e.g., marking a specific area) and deducting the next steps from both textual and visual intermediate results. The requirement, however, is difficult to accomplish for most VLMs, whether open-sourced or proprietary.
@@
III. We use least-to-most synthesis to construct the Vireo dataset of $50$k examples for fine-tuning a reasoner model. Extensive experiments illustrate that the reasoner can consistently and significantly enhance existing VLMs in a plug-and-play fashion across five VQA benchmarks [^2] All the code and data will be published later..
@@ 2.2 Reasoning and Tool Use
Reasoning is an important emergent ability of LLMs Wei et al. (2022a). With appropriate exemplars or prompts, LLMs can demonstrate “chain-of-thought” (CoT) behavior and solve problems through multi-step reasoning Wei et al. (2022b); Kojima et al. (2022). Encouraged by the observation, significant efforts have been made to improve the reasoning capabilities of LLMs. For example, Zhou et al. (2022) propose least-to-most prompting for complex reasoning; Yao et al. (2022) extend the ability of LLMs by interleaving reasoning and acting; Yao et al. (2024) introduce tree-of-thought (ToT) as an extension of CoT; and Besta et al. (2024) further extend the idea of ToT by considering a graph structure. In addition to innovations in methodology, reasoning abilities have also proven effective in various applications, such as table understanding Wang et al. (2023c), math problem solving Wei et al. (2022b), question answering Guan et al. (2024), and decision making Yao et al. (2022). Very recently, the research community begins to investigate the problem in multimodal models Qi et al. (2024); Yang et al. (2023); Wang et al. (2024). Different from existing work, we focus on data synthesis with open-sourced models, and develop a plug-an-play visual reasoner.
@@ 3.2 Least-to-Most Synthesis
We implement Questioner and Combiner by fine-tuning LLaMA-3-8B-Instruct [^6] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct. Due to limited resources, we obtain the training data by querying GPT-4 Wang et al. (2023b) using a few high-quality demonstrations as seeds following the self-instruct pipeline Wang et al. (2023b). Since the two tasks are relatively simple, querying GPT-4 only 10k times is enough to achieve satisfactory performance [^7] We will publish Questioner and Combiner models later.. More details are in Appendix B.
@@ 4.1 Models for the Answer Tool
To comprehensively illustrate the versatility of our Reasoner, we employ several VLMs with different sizes and architectures as the Answer tool, including (1) BLIP-2 Li et al. (2023a): It utilize a Q-Former module to integrate visual and textual information. We use the 2.7B version [^9] https://huggingface.co/Salesforce/blip2-opt-2.7b, which is the smallest model in our experiments. (2) InstructBLIP Dai et al. (2024): It expands BLIP-2 by incorporating instruction prompts into the Q-Former module. We use the 7B [^10] https://huggingface.co/Salesforce/instructblip-vicuna-7b and 13B [^11] https://huggingface.co/Salesforce/instructblip-vicuna-13b versions in our experiments. (3) LLaVA Liu et al. (2024a): It hinges on a basic projection layer to align image and text representations. We use the 13B version [^12] https://huggingface.co/llava-hf/llava-v1.6-vicuna-13b-hf.
@@ 4.2 Evaluation Datasets
Table 2: Evaluation results with different VLMs as the Answer tools. Size refers to the number of parameters of the language model in each VLM. We do not provide the result of LLaVA on ST-VQA because its output always includes abundant information, leading to nonsensical low scores using the official evaluation website.
@@ 4.4 Quality Assessment of Vireo
We conduct a human study to investigate the quality of our synthesized Vireo dataset. Specifically, we randomly sample $200$ examples from Vireo, and recruit $3$ graduate students majoring in NLP to manually check the correctness of the synthesized question and reasoning process for each instance. The check focuses on the correctness in three aspects: (1) Sub-Question. The sub-question must be accurately faithful to the profile of the tail node and can be resolved using the specified tool. (2) Argument. Based on the current sub-question, the annotators need to assess whether using the given argument to invoke the tool can yield the expected output. (3) Main Question. The annotators judge whether the main question covers all sub-questions and whether it can be decomposed into sub-questions in the correct order. Each example receives 3 labels from each of the three annotators on each aspect. Results show that the three evaluators consistently approve that all examples are correct in all aspects [^13] This might be surprising, but we double-checked the evaluation and confirmed the results. , suggesting the highly guaranteed quality of the data produced by the proposed synthesis approach.
@@ 5 Discussions
Although the main results affirmed the high quality of Vireo and the general benefits of the Reasoner to enhance VLMs on multiple benchmarks, we are still curious about the following research questions: (1) RQ1: How does the capability of the Reasoner vary with respect to the size of instructions for fine-tuning? (2) RQ2: Will the Reasoner generate negative impact on visual tasks other than VQA? (3) RQ3: What are the underlying causes of the Reasoner’s errors on the benchmarks?
5.1 RQ1: Regarding Influence of Data Size
@@
5.2 RQ2: Regarding Other Visual Tasks
@@
Table 3: Evaluation results on other visual tasks.
@@
5.3 RQ3: Regarding Error Analysis
@@ Appendix B Implementation Details
Appendix C Case Study
@@
Appendix D Example of Data Construction

### v2

@@ Abstract
Chuanqi Cheng1,2∗  Jian Guan2thanks:   Equal Contribution.  Wei Wu2thanks:   Corresponding Authors.  Rui Yan1 1Gaoling School of Artificial Intelligence, Renmin University of China  2Ant Group {chengchuanqi, ruiyan}@ruc.edu.cn {jianguanthu, wuwei19850318}@gmail.com
@@ 1 Introduction
In this work, we explore vision language models (VLMs) as a typical example of LMMs. Despite the rapid progress, state-of-the-art VLMs still face challenges in reasoning over visual content. As exemplified in Figure 1, intuitively, the question can be correctly solved following a “least-to-most” paradigm Zhou et al. (2022), in which the question is decomposed into a series of sub-questions, and an answer is deduced by resolving the sub-questions step by step. However, existing VLMs are inept at performing such multi-step reasoning because (1) Multi-step reasoning paths, like the one shown in Figure 1 (right), are rarely included in the training data Dai et al. (2024) [^1] A few datasets, such as A-OKVQA Schwenk et al. (2022), VCR Zellers et al. (2019), and ScienceQA Lu et al. (2022), contain rationales. However, the rationales merely provide explanations for the answers and thus differ significantly from the multi-step reasoning data we study in the work.. The VLMs have few opportunities to develop the reasoning capability from the subsequent post-training. And (2) different from reasoning over text, solving questions in a vision-language context may require manipulating the input image (e.g., marking a specific area) and deducting the next steps from both textual and visual intermediate results. The requirement, however, is difficult to accomplish for most VLMs, whether open-sourced or proprietary.
@@
III. We use least-to-most synthesis to construct the Vireo dataset of $50$k examples for fine-tuning a reasoner model [^2] We have released a larger version of Vireo comprising $500$k examples to spur research in the visual reasoning area.. Extensive experiments illustrate that the reasoner can consistently and significantly enhance existing VLMs in a plug-and-play fashion across five VQA benchmarks.
@@ 2.2 Reasoning and Tool Use
Reasoning is an important emergent ability of LLMs Wei et al. (2022a). With appropriate exemplars or prompts, LLMs can demonstrate “chain-of-thought” (CoT) behavior and solve problems through multi-step reasoning Wei et al. (2022b); Kojima et al. (2022). Encouraged by the observation, significant efforts have been made to improve the reasoning capabilities of LLMs. For example, Zhou et al. (2022) propose least-to-most prompting for complex reasoning; Yao et al. (2022) extend the ability of LLMs by interleaving reasoning and acting. Wen et al. (2024) generate code-form plans before low-level reasoning. In addition to innovations in methodology, reasoning abilities have also proven effective in various applications, such as table understanding Wang et al. (2023c), math problem solving Wei et al. (2022b), question answering Guan et al. (2024), and decision making Yao et al. (2022). Very recently, the research community begins to investigate the problem in multimodal models Qi et al. (2024); Yang et al. (2023); Wang et al. (2024). Different from existing work, we focus on data synthesis with open-sourced models, and develop a plug-an-play visual reasoner.
@@
2.3 Task Decomposition for Visual Reasoning
Various approaches have been proposed to tackle complex vision-language tasks through task decomposition and step-by-step reasoning. For example, Visual Programming Gupta and Kembhavi (2023) utilizes LLMs to perform visual reasoning by breaking down tasks into subroutines. Similarly, ViperGPT Surís et al. (2023) and CodeVQA Subramanian et al. (2023) decompose tasks to generate executable code. However, these methods heavily rely on the strong instruction-following capabilities of LLMs, making them less effective with smaller models. Furthermore, these approaches are susceptible to instability arising from prompt designs, choice and ordering of demonstrations, and LLM selection, even when employing powerful models like GPT-3 Zhao et al. (2021). Our work deviates significantly from these methods by focusing on fine-tuning VLMs on large-scale synthesized datasets, systematically enhancing the model’s capability to handle complex tasks. This approach enables us to achieve competitive performance using smaller, open-source models (e.g., 7B or 13B parameters), making our method more cost-effective, accessible, and suitable for practical deployment with fewer computational resources.
@@ 3.2 Least-to-Most Synthesis
We implement Questioner and Combiner by fine-tuning LLaMA-3-8B-Instruct [^6] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct. Due to limited resources, we obtain the training data by querying GPT-4 Wang et al. (2023b) using a few high-quality demonstrations as seeds following the self-instruct pipeline Wang et al. (2023b). Since the two tasks are relatively simple, querying GPT-4 only 10k times is enough to achieve satisfactory performance. More details are in Appendix B.
@@ 4.1 Models for the Answer Tool
To illustrate the versatility of our Reasoner, we employ several VLMs with different sizes and architectures as the Answer tool, including (1) BLIP-2 Li et al. (2023a): It utilize a Q-Former module to integrate visual and textual information. We use the 2.7B version [^8] https://huggingface.co/Salesforce/blip2-opt-2.7b, which is the smallest model in our experiments. (2) InstructBLIP Dai et al. (2024): It expands BLIP-2 by incorporating instruction prompts into the Q-Former module. We use the 7B [^9] https://huggingface.co/Salesforce/instructblip-vicuna-7b and 13B [^10] https://huggingface.co/Salesforce/instructblip-vicuna-13b versions in our experiments. (3) LLaVA Liu et al. (2024a): It hinges on a basic projection layer to align image and text representations. We use the 13B version [^11] https://huggingface.co/llava-hf/llava-v1.6-vicuna-13b-hf.
@@ 4.2 Evaluation Datasets
Table 2: Evaluation results with different VLMs as the Answer tools. Size refers to the number of parameters of the language model in each VLM. We do not provide the result of LLaVA on ST-VQA because its output always includes abundant information, leading to nonsensical low scores using the official evaluation website. We highlight the best results in bold.
4.3 Baselines
To comprehensively compare our approach with alternative task decomposition strategies, we devise a baseline termed “Tools” inspired by Gupta and Kembhavi (2023). Specifically, we employ a few-shot approach to decompose the initial problem into a sequence of tool-calling operations (e.g., PaddleOCR and GroundingDINO), which are subsequently executed. Notably, we leverage LLaMA-3-8B-Instruct Dubey et al. (2024) for decomposition, ensuring a comparable parameter count to the LLaVA-7B model used by our Reasoner.
@@ 4.5 Quality Assessment of Vireo
We conduct a human study to investigate the quality of our synthesized Vireo dataset. Specifically, we randomly sample $200$ examples from Vireo, and recruit $3$ graduate students majoring in NLP to manually check the correctness of the synthesized question and reasoning process for each instance. The check focuses on the correctness in three aspects: (1) Sub-Question. The sub-question must be accurately faithful to the profile of the tail node and can be resolved using the specified tool. (2) Argument. Based on the current sub-question, the annotators need to assess whether using the given argument to invoke the tool can yield the expected output. (3) Main Question. The annotators judge whether the main question covers all sub-questions and whether it can be decomposed into sub-questions in the correct order. Each example receives 3 labels from each of the three annotators on each aspect. Results show that the three evaluators consistently approve that all examples are correct in all aspects [^12] This might be surprising, but we double-checked the evaluation and confirmed the results. , suggesting the highly guaranteed quality of the synthesized data.
@@ 5 Discussions
Although the main results affirmed the high quality of Vireo and the general benefits of the Reasoner to enhance VLMs on multiple benchmarks, we are still curious about the following research questions: (1) RQ1: How does the capability of the Reasoner vary w.r.t. the size of instructions for fine-tuning? (2) RQ2: What is the relative impact of each integrated tool on the Reasoner’s overall capability? (3) RQ3: Will the Reasoner bring a negative impact on visual tasks other than VQA? (4) RQ4: Is the Reasoner still effective on more advanced VLMs? (5) RQ5: What are the underlying causes of the Reasoner’s errors on the benchmarks?
5.1 RQ1: Influence of Data Size
@@
Model GQA TextVQA Tally-S Tally-C
BLIP2 42.02 36.94 56.38 39.38
w/o OCR 40.26 32.30 54.89 37.89
w/o Highlight 39.34 36.30 20.45 28.66
w/o Grounding 40.44 31.98 20.89 28.84
InstBLIP-7B 53.60 43.42 74.99 57.78
w/o OCR 50.09 37.20 73.17 55.78
w/o Highlight 48.90 41.78 65.68 38.49
w/o Grounding 51.03 37.84 67.01 38.67
InstBLIP-13B 54.24 42.03 74.98 59.73
w/o OCR 50.25 36.56 73.19 55.78
w/o Highlight 48.90 41.70 57.51 25.11
w/o Grounding 51.82 37.46 58.61 25.14
LLaVA 60.65 59.22 80.28 68.65
w/o OCR 53.86 55.89 77.28 64.53
w/o Highlight 51.82 37.46 58.61 25.14
w/o Grounding 56.51 58.09 70.22 33.04
Table 3: Evaluation results of removing different tools.
@@
5.2 RQ2: Ablation Studies Regarding Tools
To gauge the significance of the diverse tools employed in our approach, we conducted a thorough ablation study, systematically investigating the impact of removing each tool on model performance. The findings, presented in Table 3, unequivocally demonstrate that the exclusion of any single tool adversely affects the model’s capabilities across various datasets. This observation underscores the multifaceted nature of visual tasks, which often necessitate the synergistic application of multiple sophisticated operations, rendering the reliance on a solitary tool insufficient.
5.3 RQ3: Extending to Broader Visual Tasks
@@
Table 4: Results on broader visual tasks beyond VQA.
@@
5.4 RQ4: Efficacy on More Advanced VLMs
Considering that more advanced VLMs such as Qwen-VL Bai et al. (2023) have huge potential to inherit the function of external tools such as perceiving bounding boxes in images, we convert Vireo into an end-to-end format by sequentially combining the input and output of each step in order to avoid the computation overhead from explicit tool invocation. Table 5 shows that the Qwen-VL model, fine-tuned on the end-to-end version of Vireo, significantly improves across all datasets. Notably, the application of our Reasoner does not cause a noticeable increase in time cost compared to the vanilla model (1.4s v.s. 1.1s per sample).
Model GQA TextVQA TallyQA-S TallyQA-C
Qwen-VL 57.95 67.26 75.38 37.92
  + Reasoner 61.73 74.12 81.50 68.40
Table 5: Evaluation results on Qwen-VL.
5.5 RQ5: Regarding Error Analysis
@@ 6 Conclusions
Acknowledgments
This work was supported by the National Natural Science Foundation of China (NSFC Grant No. 62122089), Beijing Outstanding Young Scientist Program NO. BJJWZYJH012019100020098, and Intelligent Social Governance Platform, Major Innovation & Planning Interdisciplinary Platform for the “Double-First Class” Initiative, Renmin University of China, the Fundamental Research Funds for the Central Universities, and the Research Funds of Renmin University of China.
@@ Appendix B Implementation Details
Model Size VLM-only Reasoner
BLIP2 2.7B 0.05 1.40
InstructBLIP 7B 0.19 1.70
InstructBLIP 13B 0.28 1.79
LLaVA 13B 0.57 2.08
Table 6: The average per-sample time cost (seconds) when using only the VLM and applying the Reasoner.
@@
Appendix C Additional Time Cost
While the Visual Reasoner offers significant performance gains across various tasks, it is crucial to acknowledge the computational overhead associated with its multi-step reasoning process. As each step necessitates additional inference time, the overall inference duration is extended compared to using a vanilla VLM. Table 6 presents a quantitative analysis of the average per-sample time cost using a single Nvidia A100 GPU.
Appendix D Case Study
@@
Appendix E Example of Data Construction